At a fixed temperature, Sulfuryl chloride, SO2Cl2 gas decomposes to sulfur dioxide and chlorine gas in a first-order reaction, with the rate constant of 2.90 x 10-4 s-1.

SO2Cl2 (g ) ⟶ SO2 (g) + Cl2 (g)

1.) If the initial C4H8 concentration is 0.225 M, what is the concentration after 900 s?

2.) Use the value of k to calculate the half-life of SO2Cl2 (g) in seconds.


its SO2Cl not C4H8
Answer #1

1. SOCl2 \rightarrow SO2 + Cl2

At t = 0 sec, [SOCl2] = 0.225 M = A

At t = 900 sec, [SOCl2] = 0.225-x M = At

Now,

For first order reaction-

At = A*e-kt

k= Rate constant = 2.9*10-4 s-1

t = Time = 900 sec

0.225-x = 0.225 * e(-2.9*10-4)(900) = 0.173

At = 0.173 M

Thus,

[SOCl2] after 900 sec = 0.173 M

2. Half life = 0.693/k

= 0.693 / 2.9*10-4 s-1 = 2389.66 sec = 2390 sec